OpenStudy (dls):

\[\LARGE y=10^{5\log _{10}x} \]

OpenStudy (dls):

\[\LARGE y=10^{5\log _{10}x }\]

OpenStudy (dls):

Attempt: \[\LARGE logy=5logxlog10\]

OpenStudy (dls):

\[\LARGE \frac{dy}{dx}=ylog10 \frac{5}{x}\]

hartnn (hartnn):

\(\huge a^{\log_ax}=x\)

OpenStudy (dls):

:O

hartnn (hartnn):

\(\LARGE y=10^{5\log _{10}x }=\LARGE 10^{\log _{10}x^5 }=x^5\)

hartnn (hartnn):

:O

OpenStudy (dls):

came across it first time o.o

hartnn (hartnn):

if you want to prove that, take log on both sides of that equation, and use log_p q = log q/log p

OpenStudy (dls):

I took logs on my attempt^ are u referring to that?

hartnn (hartnn):

i was saying if you want proof of that 5th ptoperty which i used here also, u take log for this Q, u directly get y=x^5

OpenStudy (dls):

\[\LARGE y=10^{5logx} \]

hartnn (hartnn):

5 log x = log (x^5)

OpenStudy (dls):

\[\LARGE 5\log _{10}xlog10\] that is what we get on taking logs

hartnn (hartnn):

why did u take log for this Q ?

OpenStudy (dls):

:o

OpenStudy (dls):

i see !